OpenStudy (anonymous):

When 5.0 mol Al react with 13 mol HCl, what is the limiting reactant, and how many moles of AlCl3 can be formed? 2 Al + 6HCl → 2 AlCl3 + 3 H2 A. Al is the limiting reactant, 5.0 mol AlCl3 can be formed B. HCl is the limiting reactant, 4.3 mol AlCl3 can be formed C. Al is the limiting reactant, 7.5 mol AlCl3 can be formed D. HCl is the limiting reactant, 6.5 mol AlCl3 can be formed

OpenStudy (anonymous):

please help<3 I'd really appreciate it

OpenStudy (mathmate):

Concentrate on the left hand side of the equation, and observe the ratio of reactants.

OpenStudy (anonymous):

Sorry I dont understand

OpenStudy (mathmate):

According to the equation of the reaction, what is the ratio of Al to HCl?

OpenStudy (anonymous):

umm 2:3 ?

OpenStudy (mathmate):

How did you arrive at 2:3?

OpenStudy (anonymous):

because the right hand side says its 2 AlCl and 3H2

OpenStudy (anonymous):

oh wait nvm it's 2:6

OpenStudy (mathmate):

The right hand side refer to products. We are dealing with reactants, meaning the substances on the left hand side, Al and HCl. Can you try again?

OpenStudy (anonymous):

Yea so it would be 2:6 or 1:4

OpenStudy (mathmate):

Yes, 2:6, can you reduce it?

OpenStudy (mathmate):

Sorry, \(2:6 \ne 1;4\)! :(

OpenStudy (anonymous):

oh lol I mean 1:3

OpenStudy (mathmate):

ok!

OpenStudy (anonymous):

ok so now what

OpenStudy (mathmate):

Now look at the quantity of actual reactants. Which of Al or HCl will be used up first if the reaction requires a ratio of 1:3?

OpenStudy (anonymous):

3 h2?

OpenStudy (anonymous):

I think I mean Im not sure

OpenStudy (mathmate):

Don't understand what you wrote! :(

OpenStudy (anonymous):

Like in the right side, theres 3 H2

OpenStudy (anonymous):

So that would be the one that'd be used up first. Idk I'm just guessing

OpenStudy (mathmate):

Yes, but do not guess, work it out!

OpenStudy (anonymous):

wait so I'm right?

OpenStudy (anonymous):

that isn't an option though

OpenStudy (mathmate):

Right now, we are not onto the right side yet. We are trying to find the limiting reagent.

OpenStudy (anonymous):

oh..ok so then I guess it would be HCl

OpenStudy (mathmate):

Limiting reagent is, the one which will be used up first.

OpenStudy (anonymous):

ohh nvm so it would be Al

OpenStudy (mathmate):

Please, do not guess. Every answer has a reason behind it. You cannot guess your way to your degree, there are exams. If you do not understand, you will have problems later on.

OpenStudy (anonymous):

No I know, I'm not guessing. I really think that it's Al

OpenStudy (mathmate):

Work it out this way. The ratio of the equation gives you the ratio of number of moles that will be used up. 1 mole of Al will require 3 moles of HCl. So how many moles of Al will use up 13 moles of HCl? Use proportions. 1 : 3 = x : 13 what is x.

OpenStudy (anonymous):

pellet man. I have no idea how to work that out. I mean, I guess it would be 1 as well? Or 10?

OpenStudy (anonymous):

I mean, it might be 10.

OpenStudy (mathmate):

Frankly, do you just want the answer, or do you want to know how to solve this problem?

OpenStudy (anonymous):

Both would be great to be honest. I just want to finish this test and get on with it. But I don't want to make it seem like I'm lazy or anything. I just need to finish it as soon as possible.

OpenStudy (mathmate):

I am sorry, but I don't see efforts in finding out how to solve the problem. Perhaps someone else will give you the answer, perhaps not. I wish you luck with your studies.

OpenStudy (anonymous):

Please help guys. Ive taken this test 4 times, I still don't understand this question, and I just need someone to assist me.

OpenStudy (mayankdevnani):

\[\large \bf 2Al+6HCl \rightarrow 2AlCl_3+3H_2\]

OpenStudy (mayankdevnani):

do you know how to find limiting reactant? @brriiarr

OpenStudy (anonymous):

no

OpenStudy (anonymous):

hello?

OpenStudy (mayankdevnani):

so here it is :- to find limiting reactant, first find stoichiometry coefficients and then mole of that particular compound and divide mole by stoichiometry coefficient and find the minimum ratio of the all the compounds and compare them and if any there is minimum ratio against all the compounds ,that compound is a limiting reactant.

OpenStudy (mayankdevnani):

For better understanding,take example of your question

OpenStudy (anonymous):

ahh ok. I sort of remember now.

OpenStudy (mayankdevnani):

ready to solve this one ? @brriiarr and if you have any problem in my step please comment. ok?

OpenStudy (anonymous):

ok yes definitely thank you so much

OpenStudy (mayankdevnani):

\[\large \bf \frac{mole}{stoichiometry~coefficient}\] \[\large \bf Al=\frac{5}{2}=2.5\] \[\large \bf HCl=\frac{13}{6}=2.16\]

OpenStudy (mayankdevnani):

which one is smaller?

OpenStudy (anonymous):

HCl

OpenStudy (anonymous):

So HCl would be one of the answers, right? Like out of the two HCl options, one is going to be correct?

OpenStudy (mayankdevnani):

correct. So HCl is limiting reactant

OpenStudy (anonymous):

ok. now what?

OpenStudy (mayankdevnani):

and to find how many moles of compound formed,we use limiting reactant and then balance it.

OpenStudy (mayankdevnani):

now i will explain it from your question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hello?

OpenStudy (mayankdevnani):

Limiting reactant is HCl,now use this :- \[\large \bf 6~moles~of~HCl \rightarrow 2~moles~of~AlCl_3\] \[\large \bf 1~mole~ HCl \rightarrow \frac{2}{6}~moles~AlCl_3=\frac{1}{3}~moles~AlCl_3\] but we have 13 moles of HCl, \[\large \bf 13~moles~of~HCl \rightarrow 13 \times \frac{1}{3}~moles~of~AlCl_3\] so our final answer, \[\large \bf 13~moles~of~HCl \rightarrow 4.3~moles~of~AlCl_3\] therefore, 4.3 moles of AlCl3 formed.

OpenStudy (mayankdevnani):

From this, option B is correct. :)

OpenStudy (anonymous):

ok thank you<3

OpenStudy (mayankdevnani):

Welcome :) <3 <3